\begin{thebibliography}{10}

\bibitem{ab}
M.~F. Atiyah and R.~Bott.
\newblock The {Y}ang-{M}ills equations over {R}iemann surfaces.
\newblock {\em Philos. Trans. Roy. Soc. London Ser. A}, 308(1505):523--615,
  1983.

\bibitem{bar}
V.~Yu. Baranovski{\u\i}.
\newblock The cohomology ring of the moduli space of stable bundles with odd
  determinant.
\newblock {\em Izv. Ross. Akad. Nauk Ser. Mat.}, 58(4):204--210, 1994.

\bibitem{bmw}
Peter Bouwknegt, Varghese Mathai, and Siye Wu.
\newblock Bundle gerbes and moduli spaces.
\newblock {\em J. Geom. Phys.}, 62(1):1--10, 2012.

\bibitem{bkvanishing}
Michel Brion and Ivan Kausz.
\newblock Vanishing of top equivariant {C}hern classes of regular embeddings.
\newblock {\em Asian J. Math.}, 9(4):489--496, 2005.

\bibitem{donaldson}
S.~K. Donaldson.
\newblock Gluing techniques in the cohomology of moduli spaces.
\newblock pages 137--170, 1993.

\bibitem{earl}
Richard Earl.
\newblock The {M}umford relations and the moduli of rank three stable bundles.
\newblock {\em Compositio Mathematica}, 109(1):13--48, 1997.

\bibitem{ek}
Richard Earl and Frances Kirwan.
\newblock The {P}ontryagin rings of moduli spaces of arbitrary rank holomorphic
  bundles over a {R}iemann surface.
\newblock {\em Journal of the London Mathematical Society}, 60:835--846, 12
  1999.

\bibitem{ek2}
Richard Earl and Frances Kirwan.
\newblock Complete sets of relations in the cohomology rings of moduli spaces
  of holomorphic bundles and parabolic bundles over a {R}iemann surface.
\newblock {\em Proc. London Math. Soc. (3)}, 89(3):570--622, 2004.

\bibitem{gieseker}
D.~Gieseker.
\newblock A degeneration of the moduli space of stable bundles.
\newblock {\em J. Differential Geom.}, 19(1):173--206, 1984.

\bibitem{salher}
R.~Herrera and S.~Salamon.
\newblock Intersection numbers on moduli spaces and symmetries of a {V}erlinde
  formula.
\newblock {\em Comm. Math. Phys.}, 188(3):521--534, 1997.

\bibitem{jkkw}
Lisa Jeffrey, Young-Hoon Kiem, Frances~C. Kirwan, and Jonathan Woolf.
\newblock Intersection pairings on singular moduli spaces of bundles over a
  {R}iemann surface and their partial desingularisations.
\newblock {\em Transform. Groups}, 11(3):439--494, 2006.

\bibitem{ljjwtoric}
Lisa Jeffrey and Jonathan Weitsman.
\newblock Toric structures on the moduli space of flat connections on a
  {R}iemann surface. {II}. {I}nductive decomposition of the moduli space.
\newblock {\em Math. Ann.}, 307(1):93--108, 1997.

\bibitem{jeffrey}
Lisa~C. Jeffrey.
\newblock Extended moduli spaces of flat connections on {R}iemann surfaces.
\newblock {\em Math. Ann.}, 298(4):667--692, 1994.

\bibitem{lisajeffrey}
Lisa~C. Jeffrey.
\newblock The {V}erlinde formula for parabolic bundles.
\newblock {\em J. London Math. Soc. (2)}, 63(3):754--768, 2001.

\bibitem{jk}
Lisa~C. Jeffrey and Frances~C. Kirwan.
\newblock Intersection pairings in moduli spaces of holomorphic bundles on a
  {R}iemann surface.
\newblock {\em Electron. Res. Announc. Amer. Math. Soc.}, 1(2):57--71
  (electronic), 1995.

\bibitem{lk}
Lisa~C. Jeffrey and Frances~C. Kirwan.
\newblock Intersection theory on moduli spaces of holomorphic bundles of
  arbitrary rank on a {R}iemann surface.
\newblock {\em Ann. of Math. (2)}, 148(1):109--196, 1998.

\bibitem{ljjw}
Lisa~C. Jeffrey and Jonathan Weitsman.
\newblock Symplectic geometry of the moduli space of flat connections on a
  {R}iemann surface: inductive decompositions and vanishing theorems.
\newblock {\em Canad. J. Math.}, 52(3):582--612, 2000.

\bibitem{kiem}
Young-Hoon Kiem.
\newblock Intersection cohomology of representation spaces of surface groups.
\newblock {\em Internat. J. Math.}, 17(2):169--182, 2006.

\bibitem{kiemli}
Young-Hoon Kiem and Jun Li.
\newblock Vanishing of the top {C}hern classes of the moduli of vector bundles.
\newblock {\em J. Differential Geom.}, 76(1):45--115, 2007.

\bibitem{kn}
A.~D. King and P.~E. Newstead.
\newblock On the cohomology ring of the moduli space of rank {$2$} vector
  bundles on a curve.
\newblock {\em Topology}, 37(2):407--418, 1998.

\bibitem{kirwan2}
Frances Kirwan.
\newblock On the homology of compactifications of moduli spaces of vector
  bundles over a {R}iemann surface.
\newblock {\em Proc. London Math. Soc. (3)}, 53(2):237--266, 1986.

\bibitem{kirwan}
Frances Kirwan.
\newblock The cohomology rings of moduli spaces of bundles over {R}iemann
  surfaces.
\newblock {\em J. Amer. Math. Soc.}, 5(4):853--906, 1992.

\bibitem{alinaquot}
Alina Marian.
\newblock On the intersection theory of {Q}uot schemes and moduli of bundles
  with sections.
\newblock {\em J. Reine Angew. Math.}, 610:13--27, 2007.

\bibitem{mocounts}
Alina Marian and Dragos Oprea.
\newblock Counts of maps to {G}rassmannians and intersections on the moduli
  space of bundles.
\newblock {\em J. Differential Geom.}, 76(1):155--175, 2007.

\bibitem{meinrenken}
E.~Meinrenken.
\newblock Witten's formulas for intersection pairings on moduli spaces of flat
  {$G$}-bundles.
\newblock {\em Adv. Math.}, 197(1):140--197, 2005.

\bibitem{nr}
M.~S. Narasimhan and S.~Ramanan.
\newblock Deformations of the moduli space of vector bundles over an algebraic
  curve.
\newblock {\em Ann. Math. (2)}, 101:391--417, 1975.

\bibitem{newstead}
P.~E. Newstead.
\newblock Characteristic classes of stable bundles of rank {$2$} over an
  algebraic curve.
\newblock {\em Trans. Amer. Math. Soc.}, 169:337--345, 1972.

\bibitem{ramanan}
S.~Ramanan.
\newblock The moduli spaces of vector bundles over an algebraic curve.
\newblock {\em Math. Ann.}, 200:69--84, 1973.

\bibitem{seshadri}
C.~S. Seshadri.
\newblock Degenerations of the moduli spaces of vector bundles on curves.
\newblock In {\em School on {A}lgebraic {G}eometry ({T}rieste, 1999)}, volume~1
  of {\em ICTP Lect. Notes}, pages 205--265. Abdus Salam Int. Cent. Theoret.
  Phys., Trieste, 2000.

\bibitem{st}
Bernd Siebert and Gang Tian.
\newblock Recursive relations for the cohomology ring of moduli spaces of
  stable bundles.
\newblock {\em Turkish J. Math.}, 19(2):131--144, 1995.

\bibitem{sikora}
Adam~S. Sikora.
\newblock Character varieties.
\newblock {\em Trans. Amer. Math. Soc.}, 364(10):5173--5208, 2012.

\bibitem{tw}
Constantin Teleman and Christopher~T. Woodward.
\newblock The index formula for the moduli of {$G$}-bundles on a curve.
\newblock {\em Ann. of Math. (2)}, 170(2):495--527, 2009.

\bibitem{thaddeus}
Michael Thaddeus.
\newblock Conformal field theory and the cohomology of the moduli space of
  stable bundles.
\newblock {\em J. Differential Geom.}, 35(1):131--149, 1992.

\bibitem{jw}
Jonathan Weitsman.
\newblock Geometry of the intersection ring of the moduli space of flat
  connections and the conjectures of {N}ewstead and {W}itten.
\newblock {\em Topology}, 37:115--132, 1998.

\bibitem{witten}
Edward Witten.
\newblock On quantum gauge theories in two dimensions.
\newblock {\em Comm. Math. Phys.}, 141(1):153--209, 1991.

\bibitem{yoshida}
Takahiko Yoshida.
\newblock The generating function for certain cohomology intersection pairings
  of the moduli space of flat connections.
\newblock {\em J. Math. Sci. Univ. Tokyo}, 8(3):541--558, 2001.

\bibitem{zaig}
Don Zagier.
\newblock On the cohomology of moduli spaces of rank two vector bundles over
  curves.
\newblock In {\em The moduli space of curves ({T}exel {I}sland, 1994)}, volume
  129 of {\em Progr. Math.}, pages 533--563. Birkh\"auser Boston, Boston, MA,
  1995.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{fullpage}
\usepackage{amsmath, amssymb, amsthm, amscd}
\usepackage{color}
\usepackage{mathrsfs}
\usepackage{amsaddr}

\newcommand{\qqq}{\mathbb{Q}}
\newcommand{\ccc}{\mathbb{C}}
\newcommand{\xxx}{X^{(2)}}
\newcommand{\wer}{{(2)}}
\newcommand{\xij}{x_{ij}}
\newcommand{\cxij}{\mathbb{Q}[x_{ij}]}
\newcommand{\cij}{\mathbb{C}_{(ij)}}

\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}

\newtheorem*{thm:main}{Theorem \ref{mainintro}}

\theoremstyle{definition}
\newtheorem{defn}[thm]{Definition}

\newtheorem*{notn}{Notation}
\newtheorem{remark}[thm]{Remark}

\theoremstyle{remark}

\title[Geometry of the intersection ring]{Geometry of the intersection ring and vanishing relations in the cohomology of the moduli space of parabolic bundles on a curve}

\author{Elisheva Adina Gamse and Jonathan Weitsman }
\address{Department of Mathematics, Northeastern University}
\thanks{Partially supported by NSF grant DMS 12--11819}
\date{\today}

\begin{document}

\begin{abstract}
We study the ring generated by the Chern classes of tautological line bundles on the moduli space of parabolic bundles of arbitrary rank on a Riemann surface. We show the Poincar\'e duals to these Chern classes have simple geometric representatives. We use this construction to show that the ring generated by these Chern classes vanishes below the dimension of the moduli space, in analogy with the Newstead-Ramanan conjecture for stable bundles.
\end{abstract}
\maketitle
\section{Introduction}

Let $G$ be a compact Lie group with maximal torus $T$, let $\Sigma^g$ be a compact, connected, oriented $2$-manifold of genus $g$, and let $p \in \Sigma^g$. For $t \in T$, let 
\[
R_g(t) = \{(A_1, \ldots, A_g, B_1, \ldots, B_g) \in G^{2g} \vert \prod_{i=1}^g [A_i, B_i] \sim t \}
\]
(where $\sim$ denotes conjugacy in $G$). The fundamental group $\pi_1(\Sigma^g \setminus \{p\})$ can be presented by generators $a_1, \ldots, a_g, b_1, \ldots, b_g, c$ with the relation $\prod_{i=1}^g [a_i, b_i]=c$, where $c$ can be thought of as representing the boundary curve of a small disc containing $p$; we choose such a set of generators. Then
\begin{equation*}
R_g(t) = \{ \rho \in Hom(\pi_1(\Sigma^g \setminus \{p\}),SU(N)) \vert \rho(c) \sim t\},
\end{equation*}
$G$ acts on $R_g(t)$ by conjugation, and $S_g(t)=R_g(t)/G$ is the space of characters of the fundamental group of $\Sigma^g \setminus \{p\}$ in $G$ where the conjugacy class of the image of $c$ is fixed.

When $t=e$, we get the moduli space $\bar{S_g} = \text{Hom}(\pi_1(\Sigma^g),G)/G$ of flat connections on $\Sigma^g$. 

Take $t=\xi \in Z(G)$ to be in the centre of $G$. In particular, take $G=SU(N)$ and $\xi = e^{2\pi i k / N}I \in Z(G)$, where $(k,N)=1$. Then the space 
\[
R_g(\xi) = \{(A_1, \ldots,A_g, B_1, \ldots, B_g) \in G^{2g} \vert \prod_{i=1}^g [A_i, B_i] = \xi\}
\]
is the moduli space of flat connections on a principal $G$-bundle over $\Sigma^g$. If we equip $\Sigma^g$ with a conformal structure, then $S_g(\xi)$ acquires a K\"ahler structure as a moduli space of stable rank $N$ vector bundles, of degree $k$ and fixed determinant, over the corresponding Riemann surface. In this case, the following generalisations of the Newstead-Ramanan conjecture have been established: 

\begin{thm}[\cite{tw}] The Chern classes of $S_g(\xi)$ vanish above degree $N(N-1)(g-1)$. 
\end{thm}
\begin{thm}[\cite{ek}] \label{genthis} The ring generated by the Chern classes of the vector bundle associated to $R_g(\xi) \to S_g(\xi)$ via the standard representation of $SU(N)$ on $\ccc^N$ vanishes in dimension strictly greater than $2N(N-1)(g-1)$. 
\end{thm}
Let $a_2, \ldots, a_r$ be the Chern classes of the vector bundle associated to $R_g(\xi)$ via the standard representation of $SU(N)$ on $\ccc^N$. Earl shows in \cite{earl} that the Pontryagin ring of $S_g(\xi)$ is contained in the subring of $H^*(S_g(\xi))$ generated by the $a_i$, and so Theorem \ref{genthis} implies 
\begin{thm}[\cite{ek}]
The Pontryagin ring of $S_g(\xi)$ vanishes in dimension strictly greater than $2N(N-1)(g-1)$.
\end{thm}
These results are generalisations of a conjecture of Newstead published in \cite{newstead}. For some other references on this subject, see \cite{ab, bar,bmw, bkvanishing,donaldson, ek2, gieseker, salher,  jkkw, ljjwtoric,jeffrey, lisajeffrey,jk, lk, ljjw,  kiem, kiemli, kn, kirwan2, kirwan,alinaquot,mocounts, meinrenken,     nr, ramanan, seshadri, st,sikora,  thaddeus, witten, yoshida, zaig}. 

We now take $t$ to be a generic element of $T$: 
\begin{defn} \label{generict}
Let $t_1, \ldots, t_N \in (0,1)$ be such that $t_1+\cdots + t_N \in \mathbb{Z}$, but the sum of any proper nonempty subset of the $t_j$ is not an integer, and let $t=\text{Diag}(e^{2\pi i t_1}, \ldots, e^{2\pi i t_N} )\in T$.
\end{defn}  Consider 
\[
R_g(t) = \{(A_1, \ldots, A_g, B_1, \ldots, B_g) \in G^{2g} \vert \prod_{i=1}^g [A_i, B_i] \sim t\}.
\]
Again, $G$ acts on $R_g(t)$ by conjugation; in this case, $S_g(t)=R_g(t)/G$  is a moduli space of rank $N$ vector bundles over $\Sigma^g$ with parabolic structure at the marked point $p$.

Consider the torus bundle $V_g(t) \to S_g(t)$ given by 
\[
\begin{split}
V_g(t) &= \{(A_1, \ldots, A_g, B_1, \ldots, B_g) \in G^{2g} \vert \prod_{i=1}^g [A_i, B_i]=t\} \\&= \{ \rho \in Hom(\pi_1(\Sigma^g\setminus \{p\}),SU(N))\vert \rho(c)=t\};
\end{split}
\]
then $S_g(t)=R_g(t)/G=V_g(t)/T$. 
For $1 \leq i, j \leq N$ with $i \neq j$, let $L_{ij}\to S_g(t)$ be the line bundle associated to $V_g(t)$ by the representation 
\begin{equation} \label{cijaction}
\begin{split}
T \times \ccc &\to \ccc\\
(\text{Diag}(e^{\sqrt{-1}\theta_1}, \ldots, e^{\sqrt{-1}\theta_N}),z)& \mapsto e^{\sqrt{-1}(\theta_i - \theta_j)}z.
\end{split}
\end{equation}
We will denote this representation of $T$ by $\ccc_{(ij)}$, and its weight by $\chi_{ij}$. Note that since $L_{ij} \otimes L_{jk} = L_{ik}$ and $L_{ij} = L_{ji}^*$, for $i,j,k$ all distinct, we have $c_1(L_{ij})+c_1(L_{jk})=c_1(L_{ik})$ and $c_1(L_{ij})=-c_1(L_{ji})$. We are interested in the subring of $H^*(S_g(t);\qqq)$ generated by the $c_1(L_{ij})$. 
When $G=SU(2)$, $V_g(t)$ is a circle bundle, and
\begin{thm}[\cite{jw}] \label{j} $c_1(L_{12})^{2g}=0$.
\end{thm}
We will sketch here the proof of Theorem \ref{j} found in \cite{jw}, as the purpose of this paper is to extend this technique to arbitrary rank. The idea is to find explicit geometric cycles Poincar\'e dual to the Chern class $c_1(V_g(t))$.  For $1 \leq i \leq g$, consider the sections $s_{A_i}$ of $L_{12}\to S_g(t)$ induced by the equivariant maps 
\begin{align*}
V_g(t) &\to \ccc_{(12)} \\
(A_1, \ldots, A_g, B_1, \ldots, B_g) &\mapsto (A_i)_{12},
\end{align*}
where the subscript $12$ denotes the $(1,2)$ matrix entry. These sections $s_{A_i}$ vanish on the cycles 
\[
D(A_i) = \{(A_1, \ldots, A_g, B_1, \ldots, B_g) \in V_g(t) \vert [A_i, t]=1\}/T.
\]
Define sections $s_{B_i}$ and cycles $D(B_i)$ similarly, and consider the intersection
\[
D:=D(A_1) \cap \cdots \cap D(A_g) \cap D(B_1) \cap \cdots \cap D(B_g).
\]
This is the image in $S_g(t)$ of the set of elements $(A_1, \ldots, A_g, B_1, \ldots, B_g) \in V_g(t)$ with the matrices $A_1, \ldots, A_g, B_1, \ldots, B_g$ all diagonal, so $\prod_{i=1}^g[A_i.B_i]=1$. Since $t\neq 1$ by Definition \ref{generict}, $D=\emptyset$, and so $c_1(L_{12})^{2g}=0$.

The purpose of this paper is to generalise this result to $G=SU(N)$. Our theorem is 
\begin{thm}
 \label{mainintro}
For $1 \leq i, j \leq N$ and $i \neq j$, let $k_{ij}$ be nonnegative integers. Then the cohomology class 
\begin{equation*}
{\prod_{\substack{1\leq i,j \leq N \\ i \neq j}}}c_1(L_{ij})^{k_{ij}} \in H^*(S_g(t);\qqq)
\end{equation*} 
vanishes whenever 
\begin{equation*}
\sum_{\substack{1\leq i,j \leq N \\ i \neq j}}k_{ij} \geq N(N-1)g - N + 2.
\end{equation*} 
\end{thm}

\begin{remark} The dimension of $S_g(t)$ is given by
\begin{align*}
\dim (S_g(t)) &= 2g\dim SU(N)-\dim T-\dim SU(N)\\
&= 2g(N^2-1) -(N-1) -(N^2-1)\\
&= 2g(N^2-1) - N^2 -N+2.
\end{align*}
Theorem \ref{mainintro} above says that monomials in the $c_1(L_{ij})$ vanish in degree $r$ for $r \geq 2gN(N-1) -2N +4$, which is well below the dimension of $S_g(t)$.
\end{remark}
The proof uses the same technique as \cite{jw} but the combinatorics of the intersections is much more complicated; we illustrate it here for the case $G=SU(3)$. 

For $1\leq i,j \leq 3$ with $i \neq j$, consider the line bundles $L_{ij} \to S_g(t)$ associated to $V_g(t)$ as above. Their Chern classes $c_{ij}=c_1(L_{ij})$ satisfy 
\begin{equation} \label{relationsthree}
\begin{split}
c_{ij}+c_{jk}+c_{ki} =0\\
c_{ij} + c_{ji} = 0.
\end{split}
\end{equation}
For $1 \leq m \leq g$, we can as before find sections $s_{A_m}^{ij}$ of $L_{ij} \to S_g(t)$ that are zero on 
\[
D_{ij}(A_m) = \{(A_1, \ldots, A_g, B_1, \ldots, B_g) \in V_g(t) \vert (A_m)_{ij}=0\}/T,
\]
and similarly sections $s^{ij}_{B_m}$ with zero locus $D_{ij}(B_m)$. The intersection 
\[
D:=\bigcap_{m=1}^g D_{12}(A_m)\cap D_{12}(B_m) \cap D_{13}(A_m) \cap D_{13}(B_m)
\]
is the image in $S_g(t)$ of the set of elements of $V_g(t)$ where $(A_m)_{12}=(A_m)_{13} = 0$ and $(B_m)_{12}=(B_m)_{13}=0$ for all $m$. Hence $(\prod_{m=1}^g[A_m,B_m])_{11}=1$, so again if $t$ is generic as in Definition \ref{generict}, $D=\emptyset$. Hence the monomial $c_{12}^{2g}c_{13}^{2g}$ vanishes, and similarly, so do $c_{23}^{2g}c_{21}^{2g}$ and $c_{31}^{2g}c_{32}^{2g}$. But any monomial in the $c_{ij}$ of degree at least $6g-1$ may be written using the relations \eqref{relationsthree} as a sum of monomials containing at least one such factor, and thus 
\begin{prop}
Let $G=SU(3)$. Then any monomial in the Chern classes of the $L_{ij}$ of degree at least $6g-1$ vanishes.
\end{prop}
The generalisation of this argument to higher rank requires more careful attention to the combinatorics and algebra of possible monomials.

\section{Combinatorial Preliminaries} \label{prelims}

We fix an integer $g \geq 2$; in the geometric application this is the genus of the $2$-manifold $\Sigma^g$. 

\begin{notn} For any set $X$, we will denote by $\xxx$ the set of 2-element subsets of $X$. \end{notn}

\begin{lem} \label{lines}
Let $X$ and $Y$ be finite sets with $\vert X \vert = n \geq 3$ and $\vert Y \vert \geq n(n-1)g-n+2$, and let $f:Y\to \xxx$ be a function. Then there exists $z \in X$ such that $\vert f^{-1}((X \setminus \{z\} ) ^{(2)}) \vert \geq (n-1)(n-2)g-n+3$.
\end{lem}
\begin{proof}
For $x \in X$, let $Q_x$ denote the set $(X\setminus \{x\})^{(2)}$. Suppose there is no $z$ for which $\vert f^{-1}(Q_z)\vert \geq (n-1)(n-2)g-n+3$. So for each $x \in X$, $\vert f^{-1}(Q_x)\vert \leq (n-1)(n-2)g-n+2$, and thus \begin{equation*}
\sum_{x \in X}\vert f^{-1}(Q_x)\vert \leq n(n-1)(n-2)g-n(n-2).
\end{equation*} But each element $\{u,v\} \in \xxx$ is contained in exactly $n-2$ of the sets $Q_x$. So $\sum_{x \in X}\vert f^{-1}(Q_x)\vert = (n-2)\vert Y \vert$. Hence \begin{align*}
\vert Y \vert &\leq \frac{n(n-1)(n-2)g - n(n-2)}{n-2} \\ &= n(n-1)g-n \\ &< \vert Y \vert.
\end{align*} This is a contradiction, so such a $z$ must exist. 
\end{proof}

\begin{defn} 
Let $X$ be a finite set. A \emph{block} in $X\times X$ is a subset $B \subset X \times X$ of the form $V \times (X \setminus V)$, where $V \subset X$ is a proper nonempty subset. If $\vert V \vert = h$ we call $B$ an \emph{$h$-by-$(\vert X \vert-h)$ block}.
\end{defn}

\begin{defn} For a finite set $X$, let $\mathcal{B}[X]$ be the set of blocks in $X \times X$. \end{defn}

\begin{notn} If $n \in \mathbb{N}$, we will write $[n]$ for the set $\{1, \ldots, n\}$. \end{notn}

\begin{remark} \label{linalg}
Suppose $A \in SU(N)$ and $1 \leq h \leq N$, let $V \subset [N]$ be a subset with $\vert V \vert = h$, and consider the block $B = V \times ([N] \setminus V) \subset [N] \times [N]$. If $A_{ij}=0$ for all $(i,j)\in B$, then there is some ordering of basis elements for which $A$ is upper block diagonal. More precisely, if $\sigma$ is a permutation of $[N]$ such that $\sigma(V)=[h]$, then the matrix $(C_{ij})=(A_{\sigma^{-1}(i)\sigma^{-1}(j)})$ has the form
$$\left(\begin{array}{ccc|ccc}
A_{\sigma^{-1}(1)\sigma^{-1}(1)} & \cdots & A_{\sigma^{-1}(1)\sigma^{-1}(h)} & 0 & \cdots & 0 \\
\vdots& \ddots& \vdots& \vdots &\ddots & \vdots\\
A_{\sigma^{-1}(N-h)\sigma^{-1}(1)} & \cdots & A_{\sigma^{-1}(N-h)\sigma^{-1}(h)} & 0 & \cdots & 0 \\
\hline
\vert & & \vert & \vert & & \vert \\
u_1 & \cdots & u_h & v_1 & \cdots & v_{N-h}\\
\vert & & \vert & \vert & & \vert \\

\end{array}\right).$$
Since $A$, and hence $C$, are unitary, the vectors $v_i$ form a basis for $\ccc^{N-h}$, and $u_i\cdot v_j=0$ for all $i,j$. Hence each of the $u_i$ must be zero, so the matrix $C$ is block diagonal. Hence $A_{ji}=0$ for all $(i,j)\in B$. Thus the condition $A_{ij}=0$ for all $(i,j) \in B$ implies that $A$ is block diagonal, with blocks of size $h$ and $N-h$, up to reordering of basis elements.
\end{remark}

\section{Algebraic Preliminaries} \label{polyring}

Let $X$ be a finite set. 

\begin{defn} \label{cxij} Let $\cxij$ be the ring $\qqq[\{x_{ij}\vert 1\leq i, j\leq \vert X \vert, i \neq j\}]$, where we adjoin variables $\xij$ for all ordered pairs $(i,j)$ with $i,j \in X$ and $i \neq j$.
\end{defn}

We will need the following lemmas. 

\begin{lem} \label{somez} 
Suppose $\vert X \vert = n \geq 3$, and let $p \in \cxij$ be a monomial of degree at least $n(n-1)g-n+2$. Then there exists some $z \in X$ such that if we factor $p$ as $p = qr$, where $q \in \qqq[\{x_{iz}, x_{zi} \vert i \in X \setminus \{z\}\}]$ and $r \in \qqq[\{\xij \vert i, j \in X \setminus \{z\}, i \neq j\}]$ are monomials, then $r$ has degree at least $(n-1)(n-2)g-n+3$. 
\end{lem}

\begin{proof}
Write $p =\lambda \prod_{i,j} \xij^{d_{ij}}$. Let $Y_{ij}$ be disjoint sets with $\vert Y_{ij}\vert = d_{ij}$, for each pair $(i,j)$ with $i, j \in X$ and $i\neq j$. Let $Y = \cup Y_{ij}$. We have \begin{equation*}
\vert Y \vert = \sum_{i,j} d_{ij} \geq n(n-1)g-n+2.
\end{equation*} 
Consider the function 
\begin{align*}
f: Y &\to \xxx\end{align*} given by \begin{equation*}f\vert_{Y_{ij}}=\{i,j\}.\end{equation*} Then \begin{equation*} 
\vert f^{-1}(\{i,j\})\vert = d_{ij}+d_{ji}
\end{equation*} for all $\{i,j\} \in \xxx$. By Lemma \ref{lines}, there exists $z \in X$ with \begin{equation*}
\vert f^{-1}((X\setminus \{z\})^\wer)\vert \geq (n-1)(n-2)g-n+3.
\end{equation*} But \begin{align*}\vert f^{-1}((X\setminus \{z\})^\wer)\vert &= \sum_{\{i,j\} \in (X\setminus \{z\})^\wer}d_{ij}+d_{ji} \\ &= \deg(r), \end{align*} and so $\deg(r) \geq (n-1)(n-2)g-n+3$.
\end{proof}

\begin{lem} \label{partition} 
Let $z \in X$ and let $w, h \in \mathbb{N}$ with $w+h=\vert X \vert - 1$. Let $\eta \in \qqq[\{x_{zi} \vert i \in X \setminus \{z\}\}]$ be a monomial of degree at least $\vert X \vert(\vert X \vert-1)g -\vert X \vert+2-2gwh$. Given a partition $X\setminus \{z\} = \{e_1, \ldots, e_h\} \sqcup \{f_1, \ldots, f_w\}$, factor $\eta$ as $\eta = \eta_h \eta_w$, with $\eta_h \in \qqq[\{x_{ze_i} \vert 1 \leq i \leq h\}]$ and $\eta_w \in \qqq[\{ x_{zf_i}\vert 1 \leq i \leq w\}]$. Then either the degree of $\eta_h$ is at least $gh(h+1)-h+1$, or the degree of $\eta_w$ is at least $gw(w+1)-w+1$. 
\end{lem}

\begin{proof}
Suppose $\deg(\eta_h)<gh(h+1)-h+1$. Then
\begin{align*}
\deg(\eta_w) &= \deg(\eta) - \deg(\eta_h)\\
&\geq \vert X \vert(\vert X \vert -1)g - \vert X \vert + 2 -2gwh - gh(h+1)+h\\
&=(w+h+1)(w+h)g-(w+h+1)+2 -2gwh -gh(h+1) + h\\
&=g(w^2 + 2hw +h^2 + w + h -2wh-h^2 -h) -w-h+1 +h\\
&=gw(w+1)-w+1. \qedhere
\end{align*}
\end{proof}

\begin{defn} \label{ringr} Let $I \subset \cxij$ be the ideal generated by $x_{ij}+x_{ji}$ and $x_{ij} + x_{jk} +x_{ki}$, for all triples of distinct elements $i,j,k \in X$. Let $R = \cxij / I$ be the quotient of $\cxij$ by this ideal. 
\end{defn} 
Note that the quotient preserves the grading by degree. If $\zeta \in \cxij$ we will write $[\zeta]$ for its image in $R$. 

\begin{lem} \label{forcerestintoline} Let $\xi \in \cxij$ be a monomial, and let $z \in X$. Then there exists a homogeneous polynomial $\eta \in \qqq[\{x_{zj}\vert j \in X \setminus \{z\}\}]$ of the same degree as $\xi$ such that $[\xi] = [\eta] \in R$. 
\end{lem}

\begin{proof}
If $\xi = \lambda \prod_{i,j}\xij^{d_{ij}}$ where $\lambda \in \qqq$, let\begin{equation*}
\eta = \lambda\prod_{i,j \neq z}(-x_{zi}+x_{zj})^{d_{ij}}\prod_{i \neq z}(-1)^{d_{iz}}x_{zi}^{d_{iz}}\prod_{j\neq z}x_{zj}^{d_{zj}}. 
\end{equation*}Then $[\xi]=[\eta]$, and the degree of each term of $\eta$ is equal to the degree of $\xi$.
\end{proof}

\begin{prop} \label{induction}Let $X$ be a finite set with $\vert X \vert \geq 2$. Let $R = \cxij / I$ as before. Let $\zeta \in \cxij$ be a monomial of degree at least $\vert X \vert ( \vert X \vert - 1)g - \vert X \vert + 2$. Then for each block $B \subset X \times X$ we can find a monomial $\psi_B$ in $\cxij$ such that 
\begin{equation*}
[\zeta] = \left[\sum_{B \in \mathcal{B}[X]}\psi_B\prod_{(i,j) \in B}x_{ij}^{2g}\right].
\end{equation*}
\end{prop}
\begin{proof}
By induction on $\vert X \vert$.

If $\vert X \vert = 2$, take $X = \{1,2\}$. The monomials in $\cxij$ of degree at least $2g$ are of the form $\lambda x_{12}^ax_{21}^b$ where $a+b\geq 2g$, $\lambda \in \qqq$. The set $\{(1,2)\}$ is a 1-by-1 block in $X \times X$. Since $[x_{12}]=[-x_{21}]$ in $R$, the class $[\lambda x_{12}^ax_{21}^b]=[\lambda(-1)^bx_{12}^{a+b-2g}x_{12}^{2g}] \in R$ is of the desired form.

Now suppose $\vert X \vert = n \geq 3$. Let $\zeta \in \cxij$ be a monomial of degree $d \geq n(n-1)g-n+2$. By Lemma \ref{somez}, there exists $z \in X$ such that if we factor $\zeta$ as $\zeta = qr$, where $q \in \qqq[\{x_{iz},x_{zi}\vert i \in X\setminus \{z\}\}]$ and $r \in \qqq[\{x_{ij}\vert i, j \in X\setminus \{z\}, i \neq j\}]$ are monomials, then the degree of $r$ is at least $(n-1)(n-2)g-n+3$. 

By the inductive hypothesis, for each block $C \subset (X \setminus \{z\})\times (X \setminus \{z\})$ we can find a monomial $\theta_C \in \qqq[\{x_{ij}\vert i, j \in X\setminus \{z\}, i \neq j\}]$ such that 
\begin{equation*}
[r] = \left[\sum_{C \in \mathcal{B}[X\setminus \{z\}]}\theta_C \prod_{(i,j) \in C}x_{ij}^{2g}\right],
\end{equation*} and so 
\begin{equation*}
[\zeta] = \left[qr\right] = \left[q \sum_{C \in \mathcal{B}[X\setminus \{z\}]}\theta_C \prod_{(i,j) \in C}x_{ij}^{2g}\right].
\end{equation*}
It suffices to show that each nonzero monomial in the sum can be written as a sum of terms having the desired form. For each $C$ with $\theta_C \neq 0$, consider 
\begin{equation*}
\left[q\theta_C \prod_{(i,j)\in C}x_{ij}^{2g}\right].
\end{equation*}
This is a monomial of degree $d\geq n(n-1)g-n+2$. Suppose the block $C$ is given by 
\[
C = \{e_1, \ldots, e_h\} \times \{f_1, \ldots, f_w\},
\]
where $h, w \geq 1$ and $X \setminus \{z\}$ is the disjoint union $X \setminus \{z\} =\{e_1, \ldots, e_h\} \sqcup \{f_1, \ldots, f_w\}$ (so $w+h=n-1$). By Lemma \ref{forcerestintoline}, we can find a homogeneous polynomial $p_1 + \cdots + p_m$, where $p_1, \ldots, p_m$ are monomials in $\qqq[\{x_{zj}\vert j \in X \setminus \{z\}\}]$, such that 
\[
\left[q \theta_C \prod_{(i,j)\in C}x_{ij}^{2g}\right]=\left[(p_1 + \cdots + p_m)\prod_{(i,j)\in C}x_{ij}^{2g}\right].
\] 
Again, it suffices to show that each monomial in the sum can be written as a sum of terms having the desired form, so consider 
\[
\left[p\prod_{(i,j) \in C}x_{ij}^{2g}\right],
\]
where $p \in \{p_1, \ldots, p_m\}$. Note that
\[
\deg(p)= d - 2gwh \geq n(n-1)g-n+2-2gwh.
\]Factor $p$ as $p=p_hp_w$ where $p_h \in \qqq[\{x_{ze_i}\vert 1 \leq i \leq h\}]$ and $p_w \in \qqq[\{x_{zf_j}\vert 1 \leq j \leq w\}]$ are monomials. By Lemma \ref{partition}, either $\deg(p_h) \geq gh(h+1)-h+1$ or $\deg(p_w) \geq gw(w+1)-w+1$; without loss of generality we assume the former. 

By the inductive hypothesis, for each block $D \subset \{e_1, \ldots, e_h, z\}\times \{e_1, \ldots, e_h, z\}$ we can find a monomial $\phi_D \in \qqq[\{x_{ij} \vert i, j \in \{e_1, \ldots, e_h,z\}, i \neq j\}]$ such that 
\[
\left[p_h\right] = \left[\sum_{D \in \mathcal{B}[\{e_1, \ldots, e_h, z\}]}\phi_D \prod_{(i,j) \in D}x_{ij}^{2g}\right],
\]
and so 
\[
\left[p \prod_{(i,j)\in C}x_{ij}^{2g}\right]=\left[ p_w \sum_{D \in \mathcal{B}[\{e_1, \ldots, e_h, z\}]}\phi_D \prod_{(i,j) \in D}x_{ij}^{2g}\prod_{(i,j) \in C}x_{ij}^{2g}\right].
\]
For each $D$, consider the monomial
\[
\left[p_w \phi_D \prod_{(i,j) \in D}x_{ij}^{2g}\prod_{(i,j) \in C}x_{ij}^{2g}\right].
\]Observe that $C \cap D = \emptyset$, and so 
\[
\left[p_w \phi_D \prod_{(i,j) \in D}x_{ij}^{2g}\prod_{(i,j) \in C}x_{ij}^{2g}\right] = \left[p_w \phi_D\prod_{(i,j)\in C \cup D}x_{ij}^{2g}\right].
 \]We may assume 
\[
D = \{e_{\sigma(1)}, \ldots, e_{\sigma(d)}\}\times \{e_{\sigma(d+1)}, \ldots, e_{\sigma(h)}, z\},
\]
for some permutation $\sigma$ of $[h]$ and $1 \leq d \leq h$, since $x_{ij}-x_{ji}\in I$. But
\[
C \cup D =\{e_1, \ldots, e_h\} \times \{f_1, \ldots, f_w\} \cup \{e_{\sigma(1)}, \ldots, e_{\sigma(d)}\} \times \{e_{\sigma(d+1)}, \ldots, e_{\sigma(h)}, z\}
\]
contains
\[
E : = \{e_{\sigma(1)}, \ldots, e_{\sigma(d)}\} \times \{e_{\sigma(d+1)}, \ldots, e_{\sigma(h)}, f_1, \ldots, f_w, z\} \in \mathcal{B}[X].
\]
So
\[
\left[p_w \phi_D \prod_{(i,j) \in D}x_{ij}^{2g}\prod_{(i,j) \in C}x_{ij}^{2g}\right] = \left[ \psi_E \prod_{(i,j) \in E}x_{ij}^{2g}\right]
\] for some $\psi_E \in \cxij$. We have shown that $[\zeta]$ has a representative in $\cxij$ that is a sum of monomials of this form, i.e.
\begin{equation*}
[\zeta] = \left[\sum_{B \in \mathcal{B}[X]}\psi_B \prod_{(i,j) \in B} x_{ij}^{2g}\right],
\end{equation*}
for some monomials $\psi_B$ in $\cxij$.
\end{proof}

\section{Proof of the main theorem}

\begin{defn} \label{sections} Suppose $u \in \{a_1, \ldots, a_g, b_1, \ldots, b_g\}$ is one of the chosen generators of $\pi_1(\Sigma^g\setminus \{p\})$, and define maps 
 \begin{align*}f^{ij}_u : V_g(t) &\to \ccc_{(ij)}\\
 \rho &\mapsto (\rho(u))_{ij}
 \end{align*} for each pair $(i.j)$ with $1 \leq i, j \leq N$ and $i \neq j$. These maps $f_u^{ij}$ are $T$-equivariant since $T$ acts on the matrix entry $(\rho(u))_{ij}$ with weight $\chi_{ij}$. These maps then induce sections 
 \begin{align*}
 s_u^{ij}: S_g(t) & \to V_g(t) \times_T \cij \\
 [\rho] &\mapsto (\rho, f_u^{ij}(\rho))
 \end{align*}
 of the line bundles $L_{ij}$. 
 
Let $D_u^{ij}$ be the image in $S_g(t)$ of the subspace $\{ \rho \in V_g(t) \vert (\rho(u))_{ij}=0\}$. Then the section $s_u^{ij}$ is nonzero on the complement of $D_u^{ij}$. 
\end{defn} 

To prove Theorem \ref{mainintro}, we will show that intersections of certain sets of these subspaces $D_u^{ij}$ are empty, and conclude that the corresponding polynomials in the Chern classes $c_1(L_{ij})$ are zero. This is the same technique that was used in \cite{jw}.

\begin{lem} \label{nowherezero}
Let $M$ be a manifold. Let $\mathcal{L}_i \to M$, $i = 1, \ldots, m$, be complex line bundles with sections $s_i:M\to \mathcal{L}_i$. If these sections have no common zeros, i.e. $s_1^{-1}(0) \cap \cdots \cap s_m^{-1}(0) = \emptyset$, then $c_1(\mathcal{L}_1) \cdots c_1(\mathcal{L}_m) = 0$. 
\end{lem}

\begin{proof} 
Consider the vector bundle $E:=\mathcal{L}_1 \oplus \cdots \oplus \mathcal{L}_m \to M$ with the section \begin{equation*}\sigma :=(s_1, \cdots , s_m) : M \to E.
\end{equation*} The section $\sigma$ is nowhere zero, so the Euler class $e(E) =0$. Since $c_m(E) = e(E)$, the top Chern class $c_m(E)=c_1(\mathcal{L}_1)\cdots c_1(\mathcal{L}_m)$ is equal to zero.
\end{proof}

\begin{lem} \label{vanishing}
Let $0<h<N$, let $B \subset [N] \times [N]$ be an $h$-by-$N-h$ block, and let 
\begin{equation*}
\zeta = \prod_{(i,j)\in B}c_1(L_{ij})^{2g} \in H^{4gh(N-h)}(S_g(t)).
\end{equation*} Then $\zeta = 0$. 
\end{lem}

\begin{proof}
Consider the sections $s_u^{ij}:S_g(t)\to L_{ij}$ (as defined in \ref{sections}), for $(i,j)\in B$ and generators $u \in \{a_1, \ldots, a_g, b_1, \ldots, b_g\}$ of $\pi_1(\Sigma^g\setminus \{p\})$. We have $(s_u^{ij})^{-1}(0) = D_u^{ij}$. By Lemma \ref{nowherezero}, it suffices to show that \begin{equation*}
D:=\bigcap_{(i,j)\in B}(D_{a_1}^{ij} \cap \cdots \cap D_{a_g}^{ij} \cap D_{b_1}^{ij} \cap \cdots \cap D_{b_g}^{ij})=\emptyset.
\end{equation*} 
By definition, $D$ is the image in $S_g(t)$ of the set of homomorphisms $\rho \in V_g(t)$ such that the $(i,j)^{th}$ entry of $\rho(u)$ is zero for all $(i,j)\in B$ and all $u \in \{a_1, \ldots, a_g, b_1, \ldots, b_g\}$. Suppose $\rho \in D$. By Remark \ref{linalg}, we can find a permutation $\sigma$ of $[N]$ such that the matrices
\[
(\Phi(u))_{ij}):=(\rho(u))_{\sigma^{-1}(i)\sigma^{-1}(j)},
\]
for $u \in \{a_1, \ldots, a_g, b_1, \ldots, b_g\}$ are all block diagonal with blocks of size $h$ and $N-h$. So $\prod_{i=1}^g[\Phi(a_i),\Phi(b_i)]$ is also block diagonal with blocks of size $h$ and $N-h$, and each block has determinant equal to $1$. Let $E \in SU(N)$ be a product of elementary matrices representing this permutation $\sigma$ of basis elements, so that 
\[
\rho(u) = E^\dagger \Phi(u)E \text{ for all }u\in\{a_1,\ldots, a_g,b_1,\ldots, b_g\}.
\]
In particular, the matrix $EtE^\dagger$ is obtained from $t \in T$ by permuting the diagonal entries. Then 
\begin{align*}
t = \rho(c) &= \prod_{i=1}^g[\rho(a_i), \rho(b_i)] \\
&=\prod_{i=1}^g[E^\dagger \Phi(a_i)E, E^\dagger \Phi(b_i)E]\\
&=\prod_{i=1}^g E^\dagger [\Phi(a_i), \Phi(b_i)]E \\
&=E^\dagger\left( \prod_{i=1}^g [\Phi(a_i),\Phi(b_i)]\right)E.
\end{align*}
Thus $E\rho(c)E^\dagger = EtE^\dagger = \prod_{i=1}^g[\Phi(a_i),\Phi(b_i)]$ is a diagonal matrix where the first $h$ diagonal entries have product equal to $1$. But this is impossible because for $\rho \in V_g(t)$ we chose $t=\rho(c)$ such that no $h$ of its diagonal entries could have product equal to $1$ (see Definition \ref{generict}). Hence the set $D$ of such $\rho$ is empty. 
\end{proof}

\begin{proof}[Proof of Theorem \ref{mainintro}] Let $X = [N]$ and consider the rings $\cxij$ and $R$ as in Definitions \ref{cxij} and \ref{ringr}. Let $J\subset H^*(S_g(t))$ be the subring generated by the $c_1(L_{ij})$ for $1 \leq i,j \leq N$ and $i \neq j$. Since $c_1(L_{ij})=-c_1(L_{ji})$ and $c_1(L_{ij})+c_1(L_{jk})=c_1(L_{ik})$, the map 
\begin{align*}
\pi:R &\twoheadrightarrow J \\
[x_{ij}]&\mapsto c_1(L_{ij})
\end{align*} defines a ring homomorphism. Consider the element
\[
\prod_{\substack{1 \leq i, j \leq N \\ i \neq j}}c_1(L_{ij})^{k_{ij}} \in J.
\] It has a representative 
\[
\left[ \prod_{\substack{1 \leq i, j \leq N \\ i \neq j}}x_{ij}^{k_{ij}} \right]
\] in $R$. Suppose 
\[
\sum_{\substack{1 \leq i,j \leq N \\ i \neq j}}k_{ij} \geq N(N-1)g-N+2.
\]Then by Proposition \ref{induction}, for each block $B \subset [N] \times [N]$ we can find a monomial $\theta_B$ in $\cxij$ such that 
\[
\left[ \prod_{\substack{1 \leq i,j \leq N \\ i \neq j}}x_{ij}^{k_{ij}} \right] = \left[ \sum_{B \in \mathcal{B}[X]}\theta_B \prod_{(i,j) \in B}x_{ij}^{2g}\right].
\]So
\begin{align*}
\prod_{\substack{1 \leq i, j \leq N \\ i \neq j}}c_1(L_{ij})^{k_{ij}} &= \pi \left(\left[ \sum_{B \in \mathcal{B}[X]}\theta_B \prod_{(i,j) \in B}x_{ij}^{2g}\right] \right) \\
&= \sum_{B \in \mathcal{B}[X]}\pi(\theta_B)\prod_{(i,j) \in B}c_1(L_{ij})^{2g},
\end{align*} which vanishes by Lemma \ref{vanishing}.
\end{proof}

\bibliographystyle{plain} 

\bibliography{/home/adina/Dropbox/Maths/Math/refs}

\end{document}